the length of a smal rectangular room is 4 units more than the width, the area of the room is 32 square units. what is the dimension of this room.
Answer by Cromlix(4381)

Width = x units
Length = x + 4 units
Area = Length x Width
32 units^2 = (x + 4)x
x^2 + 4x = 32
x^2 + 4x - 32 = 0
(x + 8)(x - 4)= 0
So, x = -8 (No answer)
x = 4
Width = 4 units
Length = 8 units
Hope this helps
:-)